OpenStudy (anonymous):

I do NOT want the answer, hOW could I find this... If point C is the midpoint of line AB, what is the coordinate of point B? C is point (1, 4) and A is point (5, 5)

hartnn (hartnn):

let co-ordinates of B be (m,n) from mid-point formula :(if u have) 1=(m+5)/2 and 4=(n+5)/2 can u find m and n now ?

hartnn (hartnn):

yup, thats correct

OpenStudy (anonymous):

Thanks(:

hartnn (hartnn):

welcome :)